OpenStudy (anonymous):

Solve and show step by step please? 1.-2/5x - 9 < 9/10 2.q + 12 - 2 (q-22) > 0

OpenStudy (austinl):

Ok, let's start with 1 shall we? :) \(\displaystyle -\frac{2}{5}x-9<\frac{9}{10}\) Now, what do you think the first step would be?

OpenStudy (austinl):

Hello? @nsnchris Any thoughts?

OpenStudy (anonymous):

no lol

OpenStudy (austinl):

Ok, for the most part, we can treat inequalities like equations. We just have to remember a few rules in doing so. \(\displaystyle-\frac{2}{5}x-9=\frac{9}{10}\) Any thoughts now? What would we do first?

OpenStudy (anonymous):

divide each side by - 2/5x

OpenStudy (austinl):

That isn't what I would do first. When you have one variable that you are solving for, you want it on one side and constants on the other. In this particular instance, we can simply add 9 to each side. \(\displaystyle -\frac{2}{5}x<\frac{9}{10}+\frac{90}{10}\) Which when simplified is, \(\displaystyle -\frac{2}{5}x<\frac{99}{10}\) Now, what would we do?

OpenStudy (anonymous):

now divide it?

OpenStudy (austinl):

Yep, and when we divide by a fraction, we have to multiply by its reciprocal. \(\displaystyle(-\frac{5}{2})\times -\frac{2}{5}x<\frac{99}{10}\times(-\frac{5}{2})\) Now, when you multiply or divide an inequality by a negative, you must flip the sign. \(\displaystyle x>\frac{99\times-5}{10\times2}\Leftrightarrow~x>-\frac{495}{20}\)

OpenStudy (austinl):

Make sense?

OpenStudy (anonymous):

kind of lol

OpenStudy (austinl):

That should be the answer, as it isn't reduce-able.

OpenStudy (anonymous):

yeah thanks